OpenStudy (anonymous):

How do you solve: 64^x = 16^x-1 (Solving exponential equations)

OpenStudy (ash2326):

64 ^x= 16 ^(x-1) 64= 4^3 16=4^2 4^(3x)=4^(2x-2) 3x=2x-2 x=-2

OpenStudy (anonymous):

thank you!